# Post-reduction inference for confidence sets of models

### v1

@@ Appendix A Proofs and derivations
[…]
@@
$f_{Q}(q)d\text{Vol}^{n-d}(q)=\frac{\exp(-a/2\sigma^{2})}{2\pi^{(n-d_{\theta})/2}}\sum_{m=0}^{\infty}\frac{1}{m!}\Bigl{(}\frac{\sqrt{2}q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma}\Bigr{)}^{m}\Gamma\Bigl{(}\frac{n-d_{\theta}+m}{2}\Bigr{)}d\text{Vol}^{n-d}(q),$ (A.1)
@@
On using Stirling’s asymptotic approximation to the gamma integral in the form $\Gamma(x+\alpha)/\Gamma(x)\simeq x^{\alpha}$ for large $x$ and $\alpha$ fixed we see that $\lim_{m\rightarrow\infty}M_{m}=0$, so that the series representation for $f_{Q}(q)d\text{Vol}^{n-d}(q)$ is convergent. On multiplying and dividing (A.1) by $\Gamma\{(n-d_{\theta})/2\}$ and approximating the resulting gamma ratios using Stirling’s approximation as above, we obtain the approximation
@@
[…]
is the surface area of the unit hypersphere $\mathcal{S}_{n-d_{\theta}-1}$ embedded in $\mathbb{R}^{n-d_{\theta}}$. This is expression (4.2).
A.2. Proof of Proposition 4.1
@@
The distributional statement is standard. Let $T:=\hat{\sigma}^{2}_{\text{mrcv}}/\sigma^{2}$ and $\nu=df_{1}+df_{2}$. The moment-generating function of a $\chi^{2}_{\nu}$ random variable is
$M(q)=(1-2q)^{-\nu/2},\quad q<\textstyle{\frac{1}{2}}.$
Thus, by Chernoff’s inequality, for any $\delta>0$ and $q<\textstyle{\frac{1}{2}}$,
$\mathop{\mathrm{pr}}(T-1>\delta)\leq e^{-q(\delta+1)}(1-2q)^{-\nu/2}.$
On substituting in the minimiser $q=(\delta+1-\nu)/2(\delta+1)$, this bound is
$\mathop{\mathrm{pr}}(T-1>\delta)\leq\exp\Bigl{(}-\frac{\delta+1+\nu}{2}\Bigr{)}\Bigl{(}\frac{\nu}{\delta+1}\Bigr{)}^{-\nu/2}.$
Similarly, for any $\delta\in[0,1)$,
$\mathop{\mathrm{pr}}(T-1<-\delta)\leq\;e^{q(\delta-1)}(1-2q)^{-\nu/2}\leq\;\exp\Bigl{(}\frac{\delta-1+\nu}{2}\Bigr{)}\Bigl{(}\frac{\nu}{1-\delta}\Bigr{)}^{-\nu/2}.$
Let
$u(\nu,\delta):=\exp\Bigl{(}-\frac{\delta+1+\nu}{2}\Bigr{)}\Bigl{(}\frac{\nu}{\delta+1}\Bigr{)}^{-\nu/2}.$
Since
$\exp\Bigl{(}\frac{\delta-1+\nu}{2}\Bigr{)}\Bigl{(}\frac{\nu}{1-\delta}\Bigr{)}^{-\nu/2}\leq u(\nu,-\delta),$
we have
$\mathop{\mathrm{pr}}(|T-1|>\delta)\leq 2\exp\Bigl{[}-\frac{1}{2}\Bigl{\{}(\delta+1+\nu)+\nu\log\Bigl{(}\frac{\nu}{\delta+1}\Bigr{)}\Bigr{\}}\Bigr{]}.$
A.4. Proof of Proposition 4.3
@@
$\hat{Y}^{(i)}=\tilde{Y}^{(i)}+[Y\;L](\hat{\Gamma}_{i}-\Gamma_{i})=\tilde{Y}^{(i)}+\delta 1_{n}$ (A.2)
@@
where the last equality follows by Lemma A.1 and because $\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}\asymp(n-d_{\theta})^{1/2}\sqrt{k}$ by orthonormality of the columns of $U$, which implies that the entries are of order $n^{-1/2}$ on average.
Lemma A.1.
Under the conditions of Propisition 4.3,
@@
$\frac{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}{\|U^{{\rm T}}\hat{Y}^{(i)}\|_{2}}\leq\frac{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}{\max\{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}-\|\delta_{u}1_{n-d_{\theta}}\|_{2},0\}},$ (A.3)
@@
It follows by a Taylor expansion for small $\|\delta_{u}1_{n-d_{\theta}}\|_{2}=O_{p}(a_{n}k^{1/2})$ in the right-hand side of (A.3) that
A.5. Proof of Theorem 4.1
@@
Consider a stochastic expansion of the standardised random variable
$T^{(i)}=U^{{\rm T}}\hat{Y}^{(i)}/(k\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}})^{1/2}.$
Let $\nu=df_{1}+df_{2}$ as given in the definition of the modified refitted cross-validation estimator (4.6). Then by Proposition 4.2 the entries of $T^{(i)}$ have a student-$t$ distribution with $\nu$ degrees of freedom. Let $T$ denote an arbitrary random variable with this same distribution. Its density function at $t$ is
@@
We will prove that the entries of the $(n-d_{\theta})$-dimensional vectors $T^{(i)}$ are independent. Suppose for a contradiction that they are not. Then either $Z^{(i)}$ must have dependent entries, or $B_{1}^{(i)}/\nu^{1/2}$ or subsequent terms must have dependent entries. Since $\hat{\sigma}^{2}\rightarrow\sigma^{2}$ as $\nu\rightarrow\infty$, in this limit, $T^{(i)}$ is mean-zero Gaussian with covariance matrix $I_{n-d_{\theta}}$, contradicting dependence in the entries of $Z^{(i)}$. Independence in the entries of $Z^{(i)}$ contradicts independence of the entries of $B_{1}^{(i)}/\nu^{1/2}$, $B_{2}^{(i)}/\nu$, etc. as the stochastic expansion (A.4) shows that these only depend on the corresponding entry of $Z^{(i)}$. That both $\nu$ and the dimension of $T^{(i)}$ are increasing with $n$ does not affect this argument.
@@
$\hat{R}_{n}\stackrel{{\scriptstyle d}}{{=}}R_{n}+O_{p}\Bigl{(}\max\Bigl{\{}k\nu^{-1/2}(n-d_{\theta})^{-1},k\nu^{-1}(n-d_{\theta})^{-1/2}\Bigr{\}}\Bigr{)}=R_{n}+O_{p}(kn^{-3/2}).$
Appendix B An alternative proof of Proposition 4.4
As noted in the proof of Proposition 4.4, the result as stated is a direct application of Polya’s theorem. An alternative derivation is via an application of the following proposition, which replaces the $o(1)$ term in the statement of Proposition 4.4. These rates may well be very conservative. On choosing $\zeta=\gamma^{2}/8$ in proposition B.1 and noting that
@@
for any $c<\sqrt{\gamma/2}$ we see that the error term in Proposition B.1 tends to zero, recovering Proposition 4.4.
Proposition B.1.
@@
We use a Riemann approximation on the bulk of the distribution for values of $v$ close to $b$, and control the truncation error for very large values of $v$, as the Riemann approximation becomes harder to control for large values of $v$. Lemmas B.1 and B.2 give
@@
Lemma B.1.
For $10<\epsilon_{n}<$ 1, introduce the events
@@
$\mathop{\mathrm{pr}}(\hat{V}\leq v)\leq\mathop{\mathrm{pr}}(V\leq v(1+\epsilon_{n}))\,\mathop{\mathrm{pr}}(\mathcal{E}_{+}(\epsilon_{n}))+\mathop{\mathrm{pr}}(\mathcal{E}_{+}^{c}(\epsilon_{n}))$ (B.1)
@@
$\mathop{\mathrm{pr}}(\hat{V}\leq v)\geq\mathop{\mathrm{pr}}(V\leq v(1-\epsilon_{n}))\,\mathop{\mathrm{pr}}(\mathcal{E}(\epsilon_{n}))$ (B.2)
By combining (B.1) and (B.2)
@@
$U(v,\bar{v},\epsilon_{n})$ $=$ $\max\Bigl{\{}\bigl{|}\mathop{\mathrm{pr}}(V\leq\bar{v})\,\mathop{\mathrm{pr}}(\mathcal{E}(\epsilon_{n}))-\mathop{\mathrm{pr}}(V\leq v)\bigr{|},$
@@
$|\displaystyle\bigl{|}\mathop{\mathrm{pr}}(V\leq\bar{v})\,\mathop{\mathrm{pr}}(\mathcal{E}(t))-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}\leq$$\;\bigl{|}\mathop{\mathrm{pr}}(V\leq\bar{v})-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}+\mathop{\mathrm{pr}}(\mathcal{E}^{c}(\epsilon_{n})),$
@@
Case 1: $T\geq 1$, $v\geq b$. Since the mode of the density of $V$ is at $b-2$, we use the following Riemann approximation to the relevant intractable integral
@@
[…]
@@
[…]
@@
=$=\frac{\{(1-\epsilon_{n})v\}^{b/2-1}}{2^{b/2}\Gamma(b/2)}\exp\left(-\frac{(1-\epsilon_{n})v}{2}\right)\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\left(1+\frac{i+1}{(1-\epsilon_{n})v}\right)^{b/2-1}\exp\left(-\frac{i}{2}\right)$
$\leq\frac{\{(1-\epsilon_{n})v\}^{b/2-1}}{2^{b/2-1}\Gamma(b/2)}\exp\left(-\frac{(1-\epsilon_{n})v}{2}\right)\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\left(1+\frac{i+1}{(1-\epsilon_{n})v}\right)^{b/2}\exp\left(-\frac{i}{2}\right).$ (B.3)
@@
in equation (B) and the partial geometric series is bounded, for large enough $b$, by
@@
Consider the multiplicative term in (B). For $b$ sufficiently large,
@@
The following bounds are used to control the truncation error in Lemma B.2.
Proposition B.2 (Laurant and Massart 2000, Lemma 1).

### v2

@@ Appendix B Proofs and derivations
$\frac{\exp(-a/2\sigma^{2})}{2(2\pi\sigma^{2})^{(n-d_{\theta})/2}}\sum_{m=0}^{\infty}\frac{1}{m!}\biggl(\frac{q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma^{2}}\biggr)^{m}\left(\int_{0}^{\infty}(r^{2})^{\tfrac{n-d_{\theta}+m}{2}-1}\exp(-r^{2}/2\sigma^{2})dr^{2}\right)d\text{Vol}^{n-d}(q).$
@@
(B.1) $f_{Q}(q)d\text{Vol}^{n-d}(q)=\frac{\exp(-a/2\sigma^{2})}{2\pi^{(n-d_{\theta})/2}}\sum_{m=0}^{\infty}\frac{1}{m!}\Bigl(\frac{\sqrt{2}q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma}\Bigr)^{m}\Gamma\Bigl(\frac{n-d_{\theta}+m}{2}\Bigr)d\text{Vol}^{n-d}(q),$
@@
On using Stirling’s asymptotic approximation to the gamma integral in the form $\Gamma(x+\alpha)/\Gamma(x)\simeq x^{\alpha}$ for large $x$ and $\alpha$ fixed we see that $\lim_{m\rightarrow\infty}M_{m}=0$, so that the series representation for $f_{Q}(q)d\text{Vol}^{n-d}(q)$ is convergent. On multiplying and dividing (B.1) by $\Gamma\{(n-d_{\theta})/2\}$ and approximating the resulting gamma ratios using Stirling’s approximation as above, we obtain the approximation
@@
$A=\frac{2\pi^{(n-d_{\theta})/2}}{\Gamma\bigl(\frac{n-d_{\theta}}{2}\bigr)}=\int_{\mathcal{S}_{n-d_{\theta}-1}}d\text{Vol}^{n-d_{\theta}-1}(q)$
is the surface area of the unit hypersphere $\mathcal{S}_{n-d_{\theta}-1}$ embedded in $\mathbb{R}^{n-d_{\theta}}$. This is expression (A.2).
B.2. Proof of Proposition 5.1
@@
For $j\in\{1,2\}$, let $e_{j}=|\hat{\mathcal{E}}^{(j)}|$ and $\mathcal{A}_{j}=\{\mathcal{E}\subseteq\{1,\ldots,p\}\colon\mathcal{E}^{*}\subseteq\mathcal{E},|\mathcal{E}|=e_{j}\}$. Write
(B.2) $\hat{\sigma}^{2}=\frac{\|\hat{M}_{1}Y^{(1)}\|^{2}+\|\hat{M}_{2}Y^{(2)}\|^{2}}{\nu},$
where $\hat{M}_{j}$ projects onto the null-space of $X^{(j)}_{\hat{\mathcal{E}}^{(-j)}}$. Consider an oracle estimator
(B.3) $\bar{\sigma}^{2}=\frac{\|\bar{M}_{1}Y^{(1)}\|^{2}+\|\bar{M}_{2}Y^{(2)}\|^{2}}{\nu},$
where $\bar{M}_{j}$ projects onto the null-space of $X^{(j)}_{\bar{E}^{(-j)}}$ for an arbitrary, non-random $\bar{E}^{(-j)}\in\mathcal{A}_{-}j$. Since the encompassing sets are deterministic and contain the true model, we know that
(B.4) $\bar{\sigma}^{2}\stackrel{{\scriptstyle d}}{{=}}\sigma^{2}\frac{\chi^{2}_{\nu}}{\nu}.$
Let $\mathbf{1}_{C}$ be the indicator of the containment event, $C=\{\mathcal{E}^{*}\subseteq\hat{\mathcal{E}}^{(1)}\cap\hat{\mathcal{E}}^{(2)}\}$. Taking the difference, applying the triangle inequality, and taking expectations gives
$\nu\,\mathbb{E}[|\hat{\sigma}^{2}-\bar{\sigma}^{2}|\mathbf{1}_{C}]\leq\mathbb{E}[|Y^{(1)T}(\bar{M}_{1}-\hat{M}_{1})Y^{(1)}|\mathbf{1}_{C}]+\mathbb{E}[|Y^{(2)T}(\bar{M}_{2}-\hat{M}_{2})Y^{(2)}|\mathbf{1}_{C}].$
On the event $C$, $\bar{M}_{j}X^{(j)}\theta=0$ and $\hat{M}_{j}X^{(j)}\theta=0$ for $j\in\{1,2\}$, thus
(B.5) $\mathbb{E}[|Y^{(j)T}(\bar{M}_{j}-\hat{M}_{j})Y^{(j)}|\mathbf{1}_{C}]=\mathbb{E}[|\varepsilon^{(j)T}(\bar{M}_{j}-\hat{M}_{j})\varepsilon^{(j)}|\mathbf{1}_{C}]\leq\mathbb{E}|\varepsilon^{(j)T}(\bar{M}_{j}-\hat{M}_{j})\varepsilon^{(j)}|.$
Since $\varepsilon^{(j)}$ is independent of $\hat{\mathcal{E}}^{(-j)}$ (the latter is a function of $Y^{(-j)}$), we will condition on $\hat{\mathcal{E}}^{(-j)}$ in the construction of a bound for the right hand side of (B.5). For a standard normal vector $Z$ and a non-random symmetric matrix $A$, $\mathbb{E}(Z^{T}AZ)=\text{tr}(A)$ and $\mathbb{E}[(Z^{T}AZ)^{2}]=\text{tr}(A)^{2}+2\text{tr}(A^{2})$, thus,
$\mathbb{E}[\varepsilon^{(j)T}(\bar{M}_{j}-\hat{M}_{j})\varepsilon^{(j)}\mid\hat{\mathcal{E}}^{(-j)}]$ $=$ $\sigma^{2}\text{tr}(\bar{M}_{j}-\hat{M}_{j})=0,$
$\mathbb{E}[(\varepsilon^{(j)T}(\bar{M}_{j}-\hat{M}_{j})\varepsilon^{(j)})^{2}\mid\hat{\mathcal{E}}^{(-j)}]$ $=$ $2\sigma^{4}\text{tr}\{(\bar{M}_{j}-\hat{M}_{j})^{2}\}=2\sigma^{4}[\text{tr}(\bar{M}_{j}^{2})+\text{tr}(\hat{M}_{j}^{2})-2\text{tr}(\bar{M}_{j}\hat{M}_{j})]\leq 4\sigma^{4}df_{-j}.$
We have used that both projection matrices have the same trace by assumption. By Cauchy-Schwarz,
$\mathbb{E}[|\varepsilon^{(j)T}(\bar{M}_{j}-\hat{M}_{j})\varepsilon^{(j)}|\mid\hat{\mathcal{E}}^{(-j)}]\leq 2\sigma^{2}\sqrt{df_{-j}}.$
Since this holds for all $\hat{\mathcal{E}}^{(-j)}$, we obtain the bound
$\mathbb{E}|\varepsilon^{(j)T}(\bar{M}_{j}-\hat{M}_{j})\varepsilon^{(j)}|\leq 2\sigma^{2}\sqrt{df_{-j}}.$
Combining these conclusions gives, by Markov’s inequality,
$\mathop{\mathrm{pr}}\left\{\left|\frac{\hat{\sigma}^{2}}{\sigma^{2}}-\frac{\bar{\sigma}^{2}}{\sigma^{2}}\right|>\delta\right\}\leq\mathop{\mathrm{pr}}\left\{\left|\frac{\hat{\sigma}^{2}}{\sigma^{2}}-\frac{\bar{\sigma}^{2}}{\sigma^{2}}\right|>\delta,C\right\}+\varepsilon_{n}\leq\frac{2}{\delta\nu}(\sqrt{df_{1}}+\sqrt{df_{2}})+\varepsilon_{n},$
where $\varepsilon_{n}:=\mathop{\mathrm{pr}}(\mathcal{E}^{*}\not\subseteq\hat{\mathcal{E}}^{(1)}\cap\hat{\mathcal{E}}^{(2)})$. If $|\hat{\mathcal{E}}^{(1)}|+|\hat{\mathcal{E}}^{(2)}|=o(n)$, then $2(\sqrt{df_{1}}+\sqrt{df_{2}})/\nu=O(n^{-1/2})$. Since $\varepsilon_{n}=O(n^{-1/2})$ by the statement of Proposition 5.2, there exists a constant $A>0$ such that
$\mathop{\mathrm{pr}}\left\{\left|\frac{\hat{\sigma}^{2}}{\sigma^{2}}-\frac{\bar{\sigma}^{2}}{\sigma^{2}}\right|>\delta\right\}\leq\left(\frac{1}{\delta}+1\right)\frac{A}{\sqrt{n}}.$
Upon setting $\delta=Mn^{-1/2}$, we obtain
$\mathop{\mathrm{pr}}\left\{\left|\frac{\hat{\sigma}^{2}}{\sigma^{2}}-\frac{\bar{\sigma}^{2}}{\sigma^{2}}\right|>Mn^{-1/2}\right\}\leq\left(\frac{\sqrt{n}}{M}+1\right)\frac{A}{\sqrt{n}}\to\frac{A}{M}\quad\text{as }n\to\infty.$
This establishes the stated order of distributional convergence.
B.4. Proof of Proposition 5.3
@@
(B.6) $\hat{Y}^{(i)}=\tilde{Y}^{(i)}+[Y\;L](\hat{\Gamma}_{i}-\Gamma_{i})=\tilde{Y}^{(i)}+\delta 1_{n}$
@@
where the last equality follows by Lemma B.1 and because $\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}\asymp(n-d_{\theta})^{1/2}\sqrt{k}$ by orthonormality of the columns of $U$, which implies that the entries are of order $n^{-1/2}$ on average.
Lemma B.1.
Under the conditions of Propisition 5.3,
@@
(B.7) $\frac{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}{\|U^{{\rm T}}\hat{Y}^{(i)}\|_{2}}\leq\frac{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}{\max\{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}-\|\delta_{u}1_{n-d_{\theta}}\|_{2},0\}},$
@@
It follows by a Taylor expansion for small $\|\delta_{u}1_{n-d_{\theta}}\|_{2}=O_{p}(a_{n}k^{1/2})$ in the right-hand side of (B.7) that
$\frac{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}{\|U^{{\rm T}}\hat{Y}^{(i)}\|_{2}}=1+O_{p}(a_{n}k^{1/2}).$
B.5. Proof of Theorem 5.1
@@
Consider
$\frac{U^{{\rm T}}\hat{Y}^{(i)}}{k^{1/2}\hat{\sigma}}$ $=\frac{U^{{\rm T}}\hat{Y}^{(i)}}{k^{1/2}\bar{\sigma}}\bigg(1+\frac{\hat{\sigma}^{2}-\bar{\sigma}^{2}}{\bar{\sigma}^{2}}\bigg)^{-1/2}$
$=:T^{(i)}(1+O_{p}(n^{-1/2})),$
where the second equality follows by Proposition 5.2, and $\bar{\sigma}^{2}$ is defined in Equation B.3. We now consider a stochastic expansion of the $T^{(i)}$ term. Let $\nu=df_{1}+df_{2}$ as in the statement of the Theorem, then by Equation B.4 in the proof of Proposition 5.2 the entries of $T^{(i)}$ are student-$t$ distributed with $\nu$ degrees of freedom. Let $T$ denote an arbitrary random variable with this same distribution. Its density function at $t$ is
@@
We will prove that the entries of the $(n-d_{\theta})$-dimensional vectors $T^{(i)}$ are independent. Suppose for a contradiction that they are not. Then either $Z^{(i)}$ must have dependent entries, or $B_{1}^{(i)}/\nu^{1/2}$ or subsequent terms must have dependent entries. Since $\hat{\sigma}^{2}\rightarrow\sigma^{2}$ as $\nu\rightarrow\infty$, in this limit, $T^{(i)}$ is mean-zero Gaussian with covariance matrix $I_{n-d_{\theta}}$, contradicting dependence in the entries of $Z^{(i)}$. Independence in the entries of $Z^{(i)}$ contradicts independence of the entries of $B_{1}^{(i)}/\nu^{1/2}$, $B_{2}^{(i)}/\nu$, etc. as the stochastic expansion (B.8) shows that these only depend on the corresponding entry of $Z^{(i)}$. That both $\nu$ and the dimension of $T^{(i)}$ are increasing with $n$ does not affect this argument.
@@
$\hat{R}_{n}$ $\stackrel{{\scriptstyle d}}{{=}}\big\{1+O_{p}(n^{-1/2})\big\}\bigr\{R_{n}+O_{p}\bigl(\max\bigl\{k\nu^{-1/2}(n-d_{\theta})^{-1},k\nu^{-1}(n-d_{\theta})^{-1/2}\bigr\}\bigr)\bigr\},$
from which the statement of Theorem 5.3 follows. ∎
Appendix C An alternative proof of Proposition 5.4
As noted in the proof of Proposition 5.4, the result as stated is a direct application of Polya’s theorem. An alternative derivation is via an application of the following proposition, which replaces the $o(1)$ term in the statement of Proposition 5.4. These rates may well be very conservative. On choosing $\zeta=\gamma^{2}/8$ in proposition C.1 and noting that
@@
for any $c<\sqrt{\gamma/2}$ we see that the error term in Proposition C.1 tends to zero, recovering Proposition 5.4.
Proposition C.1.
@@
We use a Riemann approximation on the bulk of the distribution for values of $v$ close to $b$, and control the truncation error for very large values of $v$, as the Riemann approximation becomes harder to control for large values of $v$. Lemmas C.1 and C.2 give
@@
Lemma C.1.
For $0<\epsilon_{n}<1$, introduce the events
@@
(C.1) $\mathop{\mathrm{pr}}(\hat{V}\leq v)\leq\mathop{\mathrm{pr}}(V\leq v(1+\epsilon_{n}))\,\mathop{\mathrm{pr}}(\mathcal{E}_{+}(\epsilon_{n}))+\mathop{\mathrm{pr}}(\mathcal{E}_{+}^{c}(\epsilon_{n}))$
@@
(C.2) $\mathop{\mathrm{pr}}(\hat{V}\leq v)\geq\mathop{\mathrm{pr}}(V\leq v(1-\epsilon_{n}))\,\mathop{\mathrm{pr}}(\mathcal{E}(\epsilon_{n}))$
By combining (C.1) and (C.2)
@@
$U(v,\bar{v},\epsilon_{n})\hskip-5.69046pt$ $=$ $\hskip-5.69046pt\max\Bigl\{\bigl|\mathop{\mathrm{pr}}(V\leq\bar{v})\,\mathop{\mathrm{pr}}(\mathcal{E}(\epsilon_{n}))-\mathop{\mathrm{pr}}(V\leq v)\bigr|,$
@@
$\bigl|\mathop{\mathrm{pr}}(V\leq\bar{v})\,\mathop{\mathrm{pr}}(\mathcal{E}(t))-\mathop{\mathrm{pr}}(V\leq v)\bigr|\leq$ $\;\bigl|\mathop{\mathrm{pr}}(V\leq\bar{v})-\mathop{\mathrm{pr}}(V\leq v)\bigr|+\mathop{\mathrm{pr}}(\mathcal{E}^{c}(\epsilon_{n})),$
@@
Case 1: $T\geq 1$, $v\geq b$. Since the mode of the density of $V$ is at b2b-2, we use the following Riemann approximation to the relevant intractable integral
@@
$=\int_{v}^{Tv}\frac{\lambda^{b/2}}{\Gamma(b/2)}x^{b/2-1}\exp\left(-\frac{x}{2}\right)dx$
$\leq\int_{v}^{(1+\epsilon_{n})v}\frac{1}{2^{b}\Gamma(b/2)}x^{b/2-1}\exp\left(-\frac{x}{2}\right)dx$
@@
$\leq\int^{v}_{(1-\epsilon_{n})v}\frac{1}{2^{b}\Gamma(b/2)}x^{b/2-1}\exp\left(-\frac{x}{2}\right)dx$
@@
$=\frac{\{(1-\epsilon_{n})v\}^{b/2-1}}{2^{b/2}\Gamma(b/2)}\exp\left(-\frac{(1-\epsilon_{n})v}{2}\right)\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\left(1+\frac{i+1}{(1-\epsilon_{n})v}\right)^{b/2-1}\exp\left(-\frac{i}{2}\right)$
(C.3) $\leq\frac{\{(1-\epsilon_{n})v\}^{b/2-1}}{2^{b/2-1}\Gamma(b/2)}\exp\left(-\frac{(1-\epsilon_{n})v}{2}\right)\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\left(1+\frac{i+1}{(1-\epsilon_{n})v}\right)^{b/2}\exp\left(-\frac{i}{2}\right).$
@@
in equation (C) and the partial geometric series is bounded, for large enough $b$, by
@@
Consider the multiplicative term in (C). For $b$ sufficiently large,
@@
The following bounds are used to control the truncation error in Lemma C.2.
Proposition C.2 (Laurant and Massart 2000, Lemma 1).